OpenStudy (anonymous):

Can I have some help with this please? Not quite sure how to solve it. Kareem will rent a car for the weekend. He can choose one of two plans. The first plan has an initial fee of $61.98 and costs an additional $0.07 per mile driven. The second plan has an initial fee of $53.98 and costs an additional $0.12 per mile driven. How many miles would Kareem need to drive for the two plans to cost the same?

OpenStudy (firekat97):

so lets say he drives m miles in each plan, how much would it cost him in each case?

OpenStudy (firekat97):

@DanHowell

OpenStudy (anonymous):

The first plan would be $61. 98 and the second would be $53.98, if that what you mean.

OpenStudy (firekat97):

not exactly, so for e.g. the first plan would cost him 61.98 + 0.07m do you see why?

OpenStudy (firekat97):

He pays the initial fee, and then the extra cost per mile, as he does "m" miles

OpenStudy (anonymous):

yeah, I see why, so the second would be 53.98 + the $0.12.

OpenStudy (firekat97):

yup but you forgot the m, so its 53.98 + 0.12m

OpenStudy (firekat97):

now what we can do is let the first case = the second case so that we can solve for m

OpenStudy (firekat97):

does that make sense?

OpenStudy (anonymous):

Not exactly..? Do you mean like plug them into each other?

OpenStudy (firekat97):

no I meant that because we have worked out the total costs for both plans, we can now let these costs equal each other and solve for m (because the question is asking us to find the miles driven for both plans to cost the same)

OpenStudy (anonymous):

ohhh understood, sorry. Alrighty, what first?

OpenStudy (firekat97):

no problemmmm so first write up an equation in which cost 1 = cost 2

OpenStudy (firekat97):

remember cost 1 = 61.98 + 0.07m and cost 2 = 53.98 + 0.12m

OpenStudy (anonymous):

I honestly do not have any idea how to do that, I feel like I'm making this waaaay too complicated sorry. Thanks for your help but I think I'm gonna try something else.

OpenStudy (firekat97):

wait don't worry about it, its all good! what I meant is just let 61.98 + 0.07m = 53.98 + 0.12m

OpenStudy (firekat97):

and now we can collect like terms and solve for m